# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

### v1

@@ 5 Conclusion
In this paper, we propose GraphVista, a unified framework addressing scalability and modality coordination in graph understanding. To improve scalability, GraphVista uses a lightweight hierarchical GraphRAG base to compress and store task-relevant information. For modality coordination, the planning agent routes tasks to the most suitable modality, text modality for simple property tasks or visual modality for complex local tasks. Future work will extend GraphVista to more complex graph types with labels and semantics.
@@
The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph understanding, while also presenting several potential extensions that could inspire future research directions.
@@ Appendix A Grena Benchmark
Existing graph understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
To address these gaps, we propose the Grena Benchmark, a large-scale, multi-task benchmark with step-level visual reasoning capability. It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of Vision-language models (VLMs) in graph domains.
@@ A.1 Scalability and Structural Diversity
A primary objective of Grena is to systematically evaluate the scalability graph understanding capabilities of VLMs. To this end, the benchmark includes graphs that span a wide range of scales, with node counts from 50 to 2,050. This diversity in size is essential for assessing how model performance varies with increasing graph size and structural complexity.
@@
A.2 Categorization of Graph Understanding Tasks
To provide a structured assessment of reasoning abilities, Grena Benchmark organizes tasks into two categories:
  • Simple QA Tasks. These tasks focus on querying basic properties and local structures of graphs, such as “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to store and retrieve graph information. With a highly regularized form, these tasks emphasize breadth of coverage and scalability.
  • Complex Reasoning Tasks. These tasks require models to integrate fine-grained topological structures and perform logical reasoning. For example, shortest-path problems demand identifying the optimal path from node $v_{A}$ to node $v_{B}$. Compared with retrieval tasks, these tasks probe deeper reasoning capabilities by requiring structural integration and inference.
@@ A.3 Process-Level Supervision for Visual Reasoning
Table 6: Task types in the Grena Benchmark: definition, expected output and category.
@@
Bipartite Detection Determine whether $G=(V,E)$ is bipartite (2-colorable). Output: boolean (is bipartite?), optionally a bipartition $(S,T)$.
Clique Detection Detect the existence of (or find) a clique of specified size $k$, or return maximal cliques. Output: boolean (exists $k$-clique?) or node set(s) forming a clique.
Common Third-Order Neighbors Given nodes $u,v$, compute the intersection $N^{(3)}(u)\cap N^{(3)}(v)$ where $N^{(3)}(\cdot)$ denotes nodes at distance exactly 3. Output: set of nodes.
Connectivity Check Decide whether $G$ is connected (single connected component). Output: boolean.
Connectivity Analysis Full analysis of connectivity: number of connected components, sizes of components, and component ID/size for a query node. Output: integer(s) and list(s) (component counts and sizes).
Critical Node Detection Identify articulation points (nodes whose removal increases the number of connected components) or nodes critical to connectivity. Output: set of node ids.
Cycle Detection Determine whether any cycle exists; optionally return an example cycle (sequence of nodes). Output: boolean and, if requested, a cycle path.
Edge Count Return the total number of edges $|E|$. Output: integer.
Edge Existence Given a pair $(u,v)$, decide whether edge $(u,v)\in E$. Output: boolean.
Find Connected Edges Given a node $v$ (or a small subgraph), return the set of incident edges or the edges that connect the queried substructure. Output: set of edges (pairs).
Highest-Degree Neighbor For a query node $v$, return the neighbor $w\in N(v)$ with the largest degree $\deg(w)$ (tie-breaking rule: smallest id or deterministic rule). Output: node id.
Neighbor Connections Measure connections among the neighbors of $v$: e.g., number of edges in the induced subgraph $G[N(v)]$. Output: integer (edge count) or list of edges.
Node Count Return the total number of nodes $|V|$. Output: integer.
Node Degree Given node $v$, return its degree $\deg(v)$ (or weighted degree if weights provided). Output: integer (or numeric).
Planarity Testing Determine whether $G$ is planar; optionally, produce a planar embedding if planar. Output: boolean and (optional) embedding.
Shortest Path Given nodes $u,v$, compute a shortest path (and its length). Supports unweighted or weighted variants. Output: sequence of nodes (path) and length.
Third-Order Neighbors For node $v$, return $N^{(3)}(v)$, the set of nodes at graph distance exactly 3 from $v$. Output: set of nodes.
Triangle Counting Count all triangles (3-cycles) in $G$, and/or enumerate them. Output: integer (triangle count) or list of 3-tuples (node triples).
@@ Appendix B Subgraph Extraction
To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$, formalized as:
@@
, where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$, and can be categorized as follows:
@@
, where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
@@
Appendix C Subgraph Pruning Algorithms
To maintain visual clarity and adhere to computational constraints, particularly for large graphs, the size of an extracted subgraph $G^{\prime}$ is bounded by a maximum node count, $N_{max}$. In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked to reduce its size while preserving the most topologically significant nodes. The specific pruning strategy is contingent upon the extraction method used, namely ego-centric or multi-centric, as detailed in Algorithm 1 and Algorithm 2, respectively.
@@ C.2 Multi-centric Subgraph Pruning
// Sort by priority: distance $\downarrow$, tier $\uparrow$, centrality $\uparrow$
Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),t(v),c(v))$
@@ Appendix E Experimental Settings
Appendix F Implementation Details
Key hyperparameters and framework configurations are detailed in Table 7.
@@ Appendix G Example of Visual Graph Thought Chains
Figure 6 illustrates the process for Triangle Detection. The chain begins by identifying the neighbors of a central node, then visually inspects for edges between those neighbors to confirm a 3-cycle. For Shortest Path finding, shown in Figure 7, the model emulates a breadth-first search, extending the path layer-by-layer from the source until the destination is reached. Finally, Figure 8 depicts the fundamental operation of Neighbor Retrieval, where the model identifies and highlights all nodes directly adjacent to a given target. These examples demonstrate how Visual Graph Thoughts provide explicit, verifiable reasoning grounded in visual evidence.
Appendix H Prompt and QA Templates
This section provides prompt templates for evaluating the graph understanding capabilities of Vision-Language Models (VLMs) within the GraphVista framework. The templates are categorized into Simple Property Tasks and Complex Local Tasks, reflecting the task routing mechanism of the planning agent as described in Section 3.3 of the main paper. Each template includes an example prompt, the expected model behavior based on the GraphVista architecture, and the rationale behind the chosen modality.
H.1 Simple Property Tasks: Text Modality with GraphRAG
For simple property tasks, the planning agent routes the query to the text-modality branch. This branch utilizes GraphRAG-based retrieval on the Hierarchical GraphRAG Base ($\mathcal{K}$) to efficiently extract factual information without needing visual inference.
H.2 Complex Local Tasks: Visual Modality with Visual Graph Thoughts
For tasks requiring nuanced topological understanding or multi-step reasoning, the planning agent delegates the task to the visual modality branch. This module extracts a relevant subgraph, visualizes it, and employs the Visual Graph Thoughts agent to perform grounded, step-by-step multimodal reasoning.

### v2

@@ 5 Conclusion
In this paper, we propose GraphVista, a unified framework for scalable and modality-coordinated understanding of graph structure. For scalability, it employs a hierarchical GraphRAG base to compress task-relevant information. For coordination, a planning agent routes tasks to the optimal modality, leveraging textual retrieval for properties and visual graph thoughts for topological reasoning. Future work will extend GraphVista to complex graphs with labels and semantics.
@@
The proposed GraphVista framework offers a promising approach for scalable, modality-coordinated graph structure understanding, while also presenting several potential extensions that could inspire future research directions.
@@ Appendix A Grena Benchmark
Existing graph structure understanding benchmarks suffer from three major limitations: (i) limited coverage of graph scales and topological structures; (ii) lack of task categorization grounded in fundamental graph properties and reasoning modalities; and (iii) insufficient support for modeling and evaluating multi-step visual reasoning in graph inference.
To address these gaps, we propose the Grena Benchmark, a large-scale, multi-task benchmark with step-level visual reasoning capability. It is designed to rigorously evaluate and advance the cognitive and reasoning abilities of VLMs in graph domains.
@@ A.1 Scalability and Structural Diversity
A primary objective of Grena is to systematically evaluate the scalability graph structure understanding capabilities of VLMs. To this end, the benchmark includes graphs that span a wide range of scales, with node counts from 50 to 2,050. This diversity in size is essential for assessing how model performance varies with increasing graph size and structural complexity.
@@
A.2 Categorization of Graph Structure Understanding Tasks
To provide a structured assessment of reasoning abilities across different modalities, Grena Benchmark organizes tasks into three categories:
  • Text-Modality Tasks. These tasks focus on basic statistical retrieval and accessing individual node attributes without requiring reasoning over complex topology. Examples include “What is the total number of nodes?” or “What is the degree of node $v_{A}$?” They primarily evaluate a model’s ability to interact with the Knowledge Graph ($\mathcal{K}$) for direct information retrieval.
  • Visual-Modality Tasks. These tasks involve reasoning over local graph structures and topological relationships within a limited subgraph. Examples include shortest-path problems (“Find the shortest path between nodes A and B”) or cycle detection (“Is node A part of any triangle?”). Unlike retrieval tasks, these require the model to utilize visual subgraph representations for efficient structural inference.
  • Modality-Collaborative Tasks. These tasks require integrating global graph understanding with fine-grained local structural analysis, necessitating a sequence of steps that alternate between modalities. For example, calculating the graph diameter involves identifying peripheral nodes via text modality followed by estimating pairwise distances via visual modality. These tasks evaluate the model’s ability to decompose complex queries into sequential sub-tasks.
@@ A.3 Process-Level Supervision for Visual Reasoning
Table 6: Taxonomy of tasks in the Grena Benchmark, covering 20 distinct graph structure understanding challenges.
@@
Articulation Point Detection Identify nodes whose removal increases the number of connected components. Output: List of node IDs.
Common 3rd-Order Neighbor ID Find nodes that are exactly 3 hops away from both query nodes $u$ and $v$. Output: Set of node IDs.
Connectivity Detection Determine whether the graph is fully connected. Output: Boolean.
Cycle Detection Determine whether the graph contains any cycles. Output: Boolean.
Edge Counting Return the total number of edges $|E|$. Output: Integer.
Edge Existence Checking Determine if a direct edge exists between a given pair of nodes $(u,v)$. Output: Boolean.
Connected Edge Identification Return all edges incident to a specific query node $v$. Output: Set of edges.
Graph Diameter Calculation Compute the longest shortest path between any two nodes in the graph. Output: Integer.
Highest Degree Neighbor ID For a query node $v$, identify the neighbor with the maximum degree. Output: Node ID.
Hub Node Path Finding Identify a path between two nodes that routes through high-degree hub nodes. Output: Node sequence.
Maximum Degree Node ID Identify the node(s) with the highest degree in the entire graph. Output: Node ID(s).
Maximum Clique Detection Find the size or set of nodes of the largest complete subgraph. Output: Integer (size) or set of nodes.
Neighbor Connection Analysis Count the number of edges within the induced subgraph of a node’s neighbors. Output: Integer.
Node Counting Return the total number of nodes $|V|$. Output: Integer.
Node Degree Identification Return the degree of a specific query node $v$. Output: Integer.
Planarity Testing Determine if the graph can be embedded in a plane without edge crossings. Output: Boolean.
Shortest Path Finding Compute the shortest path distance and sequence between nodes $u$ and $v$. Output: Path length and sequence.
Star Structure Identification Determine if a specific subgraph forms a star topology centered at node $v$. Output: Boolean.
Third-Order Neighbor ID Return the set of nodes at exactly distance 3 from a query node $v$. Output: Set of node IDs.
Triangle Counting Count the total number of triangles (3-cycles) in the graph. Output: Integer.
@@ Appendix B Subgraph Extraction
To mitigate the scalability challenges of large graphs, we first extract a task-relevant local subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ from $G$ stored in $\mathcal{K}$, based on the key entities $\mathcal{E}$. This process is formalized as:
@@
where the extraction strategy $f_{\text{extract}}$ is determined by the structural characteristics of $\mathcal{E}$ and can be categorized as follows.
@@
where $\text{dist}_{G}(u,v)$ denotes the shortest path distance between nodes $u$ and $v$ in $G$. The edge set $E^{\prime}$ is induced from $G$ by $V^{\prime}$.
@@
Appendix C Subgraph Pruning and Visualization Strategy
To maintain visual clarity and adhere to computational constraints, particularly for large graphs, the size of an extracted subgraph $G^{\prime}$ is bounded by a maximum node count, $N_{max}$. In instances where the initial extracted subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ exceeds this threshold, a principled pruning procedure is invoked. Furthermore, the remaining nodes are visualized using a task-driven strategy to ensure optimal perception by the VLM.
@@ C.2 Multi-centric Subgraph Pruning
C.3 Task-Driven Subgraph Visualization Strategy
The visual representation of the extracted subgraph $G^{\prime}$ plays a critical role in the VLM’s reasoning capability. We employ a task-driven visualization strategy where the graph layout algorithm is dynamically selected based on the task type $T$:
  • Hierarchical Layouts: For sequential reasoning tasks such as Shortest Path or Cycle Detection, we utilize hierarchical layouts (e.g., Sugiyama algorithm). This layout emphasizes directionality and flow, enabling the VLM to visually trace paths layer-by-layer effectively.
  • Force-Directed Layouts: For structural analysis tasks such as Community Detection or Clustering Coefficient, we employ force-directed algorithms (e.g., Fruchterman-Reingold). This approach naturally clusters densely connected nodes, making community structures visually distinct.
Reproducibility and Clarity.
To ensure strict reproducibility of the visual inputs, we fix the random seeds for all layout generation processes. Furthermore, to handle potential visual clutter in high-density subgraphs, we apply a visual anti-overlap mechanism. Following the pruning process, we dynamically adjust node repelling forces in the layout engine to ensure that no two nodes overlap and that edge crossings are minimized. This provides a clear “what-you-see-is-what-you-get” input for the Visual Graph Thoughts agent.
@@ C.3 Task-Driven Subgraph Visualization Strategy
// Sort priority for removal: Farthest, Tier 3 (Peripheral), Lowest Centrality
Sort $V_{\text{cand}}$ into list $L$ by key $(-d(v,v_{i}),-t(v),c(v))$
@@ Appendix E Experimental Settings
E.1 Experimental Settings and Hyperparameters
Our experiments are conducted using 8 NVIDIA H20 GPUs. The framework is implemented in Python 3.11 with CUDA Version 12.6, leveraging the VLLM library for efficient inference of the VLMs. Key hyperparameters and framework configurations are detailed in Table 7.
E.2 Process-Level DPO Training Details
We fine-tuned $\mathcal{M}_{VRA}$ (instantiated with Gemma-3-12B and Qwen2.5-VL-7B) using the curated preference dataset $\mathcal{D}$. The DPO training was performed with a regularization parameter $\beta=0.1$. We utilized a learning rate of $5e^{-7}$ with a cosine decay scheduler and an effective batch size of 16 (achieved via gradient accumulation). To ensure memory efficiency during training, we employed LoRA (Low-Rank Adaptation) with rank $r=64$ and $\alpha=128$.
The preference dataset $\mathcal{D}$ consists of 10,221 samples, specifically constructed to cover diverse reasoning failure modes. The distribution of error types in the rejected trajectories ($y_{l}$) is summarized in Table 9.
Error Type Proportion Count
Multimodal Errors 35% 3,577
Textual and Logical Errors 65% 6,644
Table 9: Distribution of error types in the rejected trajectories of the preference dataset.
This balanced distribution ensures the model learns to robustly ground its reasoning in visual data while maintaining logical coherence.
@@ E.2 Process-Level DPO Training Details
Appendix F Stability Analysis of DPO Training
To further verify the robustness of the Visual Reasoning Agent ($\mathcal{M}_{VRA}$) within GraphVista, we conducted an in-depth analysis of the training stability of Process-level Direct Preference Optimization (DPO). This section presents detailed training metrics for Qwen2.5-VL-7B fine-tuned using DPO.
F.1 Training Loss and Reward Curves
Figure 9 illustrates the evolution of loss and implicit rewards over 7,500 training steps.
(a) Training Loss
(b) Implicit Rewards
Figure 9: DPO training dynamics of Qwen2.5-VL-7B on the Grena preference dataset. The left panel shows rapid convergence of training loss; the right panel demonstrates that the model successfully widened the reward margin between Chosen and Rejected paths.
Analysis of Training Dynamics
The training process exhibited high stability without the divergence often observed in RL fine-tuning:
  • Loss Convergence: The training loss (Figure 9(a)) displays a smooth downward trend, entering a stable convergence phase around 4,500 steps. This indicates that the model effectively learned the distribution of the preference data.
  • Reward Margin Separation: As shown in Figure 9(b), the reward values for Chosen paths (indicated by the blue solid line) remain relatively stable, while the reward values for Rejected paths (indicated by the orange dashed line) decrease significantly with training steps. This indicates that the optimizer maximizes the preference margin primarily by suppressing negative samples (i.e., reasoning chains containing visual hallucinations or logical errors), aligning with the theoretical expectations of DPO.
The observed stability is primarily attributed to our mild regularization strategy ($\beta=0.1$) and the high-quality process preference dataset derived from real topological structures.
@@ Appendix G Example of Visual Graph Thought Chains
Figure 6 illustrates the process for Triangle Detection. The chain begins by identifying the neighbors of a central node, then visually inspects for edges between those neighbors to confirm a 3-cycle. For Shortest Path finding, shown in Figure 7, the model emulates a breadth-first search, extending the path layer-by-layer from the source until the destination is reached. Finally, Figure 8 depicts the fundamental operation of Neighbor Retrieval, where the model identifies and highlights all nodes directly adjacent to a given target.
Task Categorization Examples 1. Text-Modality Tasks (Attribute, Statistic, Retrieval) Query: “What is the weight of the edge between Node 5 and 9?” Logic: Targets discrete attributes. Visual parsing is redundant; direct GraphRAG retrieval is optimal.   2. Visual-Modality Tasks (Topological Reasoning) Query: “Identify a cycle of length 3 involving Node A.” Logic: Requires tracing paths. Visual inspection exploits spatial recognition, superior to textual adjacency processing.   3. Collaborative Tasks (Global-Local Composition) Query: “Find the highest degree node and list its 2-hop neighbors.” Logic: Step 1 (Text): Global search for target node. Step 2 (Visual): Local subgraph reasoning.
Figure 10: Categorization guidelines used by the Planning Agent to route tasks based on structural dependency profiles.
Appendix H Task Parsing and Routing Mechanism
This section delineates the decision-making logic of the Planning Agent, which serves as the semantic router within the GraphVista framework. Building upon the architectural overview, we detail the criteria and operational rationale for allocating tasks across Text, Visual, and Collaborative modalities.
To balance computational efficiency with semantic flexibility, the Planning Agent employs a hierarchical two-stage routing protocol:
  1. 1.
    Deterministic Template Routing (Primary Strategy): The agent initially queries a library of predefined task templates (e.g., “What is the degree of node <<ID>>?”). Upon identifying an exact match, the task is deterministically typed and immediately routed to the designated modality, bypassing complex inference to minimize latency for standardized queries.
  2. 2.
    Semantic Inference Fallback (Secondary Strategy): In the absence of a template match, the agent resorts to semantic parsing. This mechanism analyzes the structural dependency of the query to infer user intent, specifically determining whether the solution necessitates topological traversal (Visual), attribute retrieval (Text), or a synergistic combination of both (Collaborative).
Rationale for Text-Modality Assignment.
The Text Modality is explicitly designated for tasks akin to database retrieval. This category encompasses: (1) Existence Verification, such as confirming the presence of nodes or edges; (2) Attribute Retrieval, which involves extracting specific properties (e.g., weights, labels) from the knowledge base $\mathcal{K}$; and (3) Global Aggregation, for operations like degree counting. For these tasks, direct retrieval offers computational superiority over visual processing, effectively mitigating latency and eliminating the noise associated with rendering dense graphs for simple factual extraction.
Rationale for Visual-Modality Assignment.
The Visual Modality is reserved for tasks necessitating "topological perception." Large Language Models (LLMs) processing linearized graph descriptions (e.g., adjacency lists) are prone to contextual attenuation ("lost-in-the-middle" phenomena) when tracing extended paths. Conversely, the visual modality projects the graph into a 2D layout where topological features—such as clusters, bridges, and cycles—become explicitly salient. This is critical for tasks like Shortest Path or Motif Detection, where the VLM can perceive connectivity patterns more effectively than through iterative textual deduction.
Rationale for Collaborative-Modality Assignment.
Complex graph reasoning often demands a hybrid "Search-then-Reason" paradigm. Collaborative tasks are characterized by a global search space (optimally handled by GraphRAG indexing) coupled with local structural verification (optimally handled by Visual Graph Thoughts). The Planning Agent decomposes such queries into sequential sub-goals. For instance, in diameter estimation, the Text branch first filters peripheral node candidates to reduce the search space, after which the Visual branch executes fine-grained pathfinding on the induced subgraph.
H.1 Analysis of Task Classification Accuracy
To evaluate the reliability of the routing mechanism, we quantify the task classification error rates across different VLM backbones. Table 10 presents the category misclassification rates for the Planning Agent.
VLM Error Rate (%)
GLM-4.1V-9B 0.1650
InternVL3-9B 0.1593
Qwen2.5-VL-7B 0.2482
Gemma-3-12B 0.1697
Qwen2.5-VL-7B (DPO) 0.2495
Gemma-3-12B (DPO) 0.1713
Table 10: Task Classification Error Rates across different VLM backbones. Models fine-tuned with DPO are denoted with (DPO).
The results highlight two key observations regarding the routing stability:
  • Impact of Model Scale: There is a clear correlation between model capacity and routing accuracy. Larger models, such as InternVL3-9B and GLM-4.1V-9B, consistently achieve lower error rates (approx. 0.16%) compared to the 7B-parameter baselines (approx. 0.25%). This suggests that the semantic nuance required for accurate intent classification benefits significantly from stronger foundational language understanding.
  • Robustness to DPO Fine-tuning: The application of Direct Preference Optimization (DPO)—while crucial for enhancing the step-by-step visual reasoning detailed in the main methodology—does not degrade the semantic routing capability. The error rates for DPO-tuned models (e.g., Qwen2.5-VL-7B (DPO) at 0.2495) remain comparable to their base counterparts (0.2482). This indicates that our alignment strategy effectively isolates visual reasoning improvements without incurring an “alignment tax” on the model’s general instruction-following and planning abilities.
System Prompt for Planning Agent Task Parsing and Routing Role Definition: You are an expert Graph Analysis Planning Agent. Your objective is to parse user queries regarding graph structures, extract key entities, classify the specific task type, and route the task to the optimal processing modality. Parsing Logic: 1. Template Verification: Initially, verify if the user input aligns with any predefined templates (e.g., Degree, Weight, Shortest Path). If a match is detected, output the standard JSON immediately. 2. Semantic Fallback: Upon template mismatch, analyze the "structural dependency" of the request to determine the required modality (Text vs. Visual vs. Collaborative). Task Categorization Guidelines (for Fallback): Text-Modality Tasks: Tasks requiring direct retrieval of statistical data, node attributes, or global counts devoid of complex structural reasoning. Examples: Node Count, Node Degree, Edge Existence, Edge Weight. Visual-Modality Tasks: Tasks necessitating topological understanding, pathfinding, or pattern recognition within a local subgraph. Examples: Shortest Path, Cycle/Triangle Detection, Common Neighbors, Clique Detection. Modality-Collaborative Tasks: Tasks requiring global structural understanding necessitating decomposition into sequential text and visual sub-tasks. Examples: Graph Diameter, Critical Node Detection, Connectivity Analysis. Output Schema (JSON):
  "task_type": "String (e.g., shortest_path, node_degree)",
  "entities": ["List of key node IDs or null"],
  "modality": "One of [Text, Visual, Collaborative]",
  "reasoning": "Brief explanation of the modality choice",
  "decomposition": ["List of sub-steps if modality is Collaborative, else null"]
Few-Shot Demonstrations: User Input: "What is the degree of node 15?" Agent Output:
  "task_type": "node_degree", "entities": ["15"], "modality": "Text",
  "reasoning": "Template Match: ’node_degree’. Retrieves attribute via GraphRAG.",
  "decomposition": null
User Input: "Find the shortest path between node A and node B." Agent Output:
  "task_type": "shortest_path", "entities": ["A", "B"], "modality": "Visual",
  "reasoning": "Template Match: ’shortest_path’. Topology required.",
  "decomposition": null
User Input: "Analyze the network to find bottlenecks impacting flow." Agent Output:
  "task_type": "bottleneck_detection", "entities": [], "modality": "Collaborative",
  "reasoning": "Fallback: Unmatched template. Semantic parsing indicates global
  search followed by local verification.",
  "decomposition": [
      "Identify high-betweenness candidates (Text)",
      "Analyze local connectivity for flow restriction (Visual)"
Current Request: <USER_QUERY>
Figure 11: The system prompt utilized by the Planning Agent. The mechanism prioritizes template matching for efficiency, falling back to semantic parsing for complex or unstructured queries.
Appendix I Prompt and QA Templates
This section delineates the standardized prompt templates and Question-Answering (QA) protocols utilized to rigorously evaluate the graph structure understanding capabilities of Vision-Language Models (VLMs) within the proposed GraphVista framework. To ensure a comprehensive assessment of both retrieval accuracy and reasoning depth, we categorize these templates into Text-Modality Tasks and Visual-Modality Tasks. This categorization strictly adheres to the task routing logic executed by the Planning Agent, as detailed in Section 3.3 of the main paper.
I.1 Text-Modality Tasks: Retrieval via GraphRAG
The Text-Modality category encompasses tasks necessitating precise information retrieval and statistical aggregation rather than abstract topological reasoning. Upon classifying a query as text-dependent, the Planning Agent routes the task to the text-modality branch. This branch leverages the Hierarchical GraphRAG Base ($\mathcal{K}$) to execute targeted retrieval operations. By accessing the structured textual descriptions stored within $\mathcal{K}$ (e.g., node attributes, edge weights, and local neighborhoods), the model extracts factual information directly, thereby bypassing the potential ambiguity associated with visual inference for non-spatial queries.
I.2 Visual-Modality Tasks: Reasoning with Visual Graph Thoughts
This category targets tasks demanding intricate topological perception and multi-step logical inference, such as pathfinding or cycle detection. For such queries, the Planning Agent delegates execution to the visual-modality branch. This module orchestrates a pipeline consisting of: (1) dynamic subgraph extraction from $\mathcal{K}$ to isolate the region of interest; (2) task-driven visualization to render the subgraph into a high-resolution image; and (3) the deployment of the Visual Graph Thoughts Agent. This agent employs a chain-of-thought mechanism grounded in the generated visual evidence to perform iterative, step-by-step reasoning, ensuring that the final answer is topologically consistent.